# Combining Flow Matching and Transformers for Efficient Solution of Bayesian Inverse Problems

### v1

@@ 7 Conclusions
We believe that our method quite universal and can be adapted to a large number of problems in a short time if the problem is reduced to the standard Bayesian Inverse Problem formulation, since it can learn complex nonlinear distributions. Another advantage is the possibility of using an input that is not fixed in terms of the number of observations, where increasing the number of observed points increases the accuracy of the method in terms of recovering the solution from the generated parameters. Finally, we can use the learned distribution to do Bayesian optimal experiment design.
Appendix A Transformer architecture for two cases
@@
Appendix B Alternative algorithm
Input: Dataset of paired samples $(x_{1},e,d)$, neural network model $\mathbf{v}_{\theta}(t,x,e,d)$, conditioning data $e$ and $d$, time $t\sim\text{Uniform}(0,1)$, number of epochs $N_{\text{epoch}}$
Output: Trained conditional flow model $\mathbf{v}_{\theta}(t,x,e,d)$.
@@ Appendix B Alternative algorithm
      for each minibatch of samples $(x_{0},x_{1})$ do
@@
            $x_{0}\sim\text{prior distribution}$
            $x_{t}\leftarrow t\cdot x_{1}+(1-t)\cdot x_{0}$
            Compute the target velocity: $u_{t}\leftarrow x_{1}-x_{0}$
            Predict the velocity: $v_{t}\leftarrow\mathbf{v}(t,x_{t},e,d)$
            Compute the loss: $\mathcal{L}(\theta)\leftarrow\mathbb{E}\left[(v_{t}-u_{t})^{2}\right]$
@@
            Accumulate gradients: $\nabla_{\Sigma}\mathcal{L}(\theta)=\nabla_{\Sigma}\mathcal{L}(\theta)+\nabla_{\theta}\mathcal{L}(\theta)$
@@
      Update $\theta$ using the optimizer and $\nabla_{\Sigma}\mathcal{L}(\theta)$
@@
Algorithm 3 Conditional Flow Matching Training Algorithm v2

### v2

@@ 8 Conclusions
We believe that our method is quite universal and can be adapted to a large number of problems in a short time when the problem is reduced to a standard Bayesian inverse problem formulation, since it can learn complex nonlinear distributions. Another advantage is the possibility of using an input that is not fixed in terms of the number of observations, where increasing the number of observed points improves accuracy in recovering the solution from the generated parameters. Finally, we can use the learned distribution to do Bayesian optimal experiment design.
Appendix A Numerical Experiments Problem Statements
A.1 Simple nonlinear model
$d(e,m)=e^{2}m^{3}+m\exp\left(-\left|0.2-e\right|\right)+\eta$
  • Sampling $m$ from $\mathcal{U}[0,1]$
  • Sampling $e$ from $\mathcal{U}[0,1]$
  • Sampling noise $\eta$ from the noise distribution
  • Computing $d=f(m,e)+\eta$
A.2 SEIR disease model
$\frac{dS}{dt}$ $=-\beta(t)SI,\frac{dE}{dt}=\beta(t)SI-\alpha E$
$\frac{dI}{dt}$ $=\alpha E-\gamma(t)I,\frac{dR}{dt}=\gamma(t)I$
where $S(t)$, $E(t)$, $I(t)$, $R(t)$ denote the fractions of susceptible, exposed, infected, and removed individuals at time $t$, respectively. These are initialized with $S(0)=99$, $E(0)=1$, $I(0)=R(0)=0$.
$\beta(t)$ $=\beta_{1}+\frac{\tanh(7(t-\tau))}{2}(\beta_{2}-\beta_{1})$
$\gamma(t)$ $=\gamma^{r}+\gamma^{d}(t)$
We fix $\tau=2.1$ over a time interval of $[0,4]$. The experiment consists of choosing four time points $e=[a_{1},a_{2},a_{3},a_{4}]\sim\mathcal{U}[1,3]$ to measure the number of infected and deceased individuals $d_{i}=[I_{e_{i}},R_{e_{i}}]$ for $i\in[1,4]$ ($d\in\mathbf{R}^{2\times 4}$). The goal is to optimally infer the 6 rates $\mathbf{m}=[\beta_{1},\alpha,\gamma^{r},\gamma^{d}_{1},\beta_{2},\gamma^{d}_{2}]$. After training an MLP and solving the flow matching problem, we learn a smooth transition from the distribution $\mathcal{U}[0,1]^{6}$ to the distribution $\hat{\mathbf{m}}\sim\rho(\mathbf{m}|\mathbf{e},\mathbf{d})$.
To summarize the inputs and outputs:
$\varepsilon=\frac{\|d-\hat{d}\|_{2}}{\|d\|_{2}}$
A.3 Permeability field inversion
$-\nabla\cdot(\kappa\nabla u)=0$
with boundary conditions:
$m(x,\mathbf{m})\approx\sum_{i=1}^{n_{m}}m_{i}\sqrt{\lambda_{i}}\phi_{i}(x),$
The transformer architecture accommodates various input formats for this inverse problem. Here, in addition to the observed solution values, we use the coordinates of measurement points. The specific architecture is detailed in Figure 6 in Appendix B.
Appendix B Technical details
@@
Table 4: Hyperparameters for SEIR and Permeability Inversion tasks
Parameter SEIR Permeability Inversion
learning_rate 8e-4 3e-4
n_emb 32 32
n_head 4 4
n_layer 6 4
Appendix C Algorithms
C.1 Conditional Flow Matching
This section provides pseudocode for the core training and inference procedures used in our Conditional Flow Matching (CFM) framework. These algorithms form the backbone of our method for solving inverse problems in various scientific settings.
Algorithm 1 details the training procedure for the conditional flow model. Given a dataset of paired samples and conditioning information, the model is trained to approximate the velocity field that defines an interpolation between prior and posterior samples. The training objective minimizes the squared error between the predicted velocity and the ground-truth velocity vector defined by the linear interpolation between samples.
Input: Dataset of paired samples $(m_{1},e,d)$, neural network model $\mathbf{v}_{\theta}(t,m,e,d)$, conditioning data $e$ and $d$, time $t\sim\text{Uniform}(0,1)$, number of epochs $N_{\text{epoch}}$
Output: Trained conditional flow model $\mathbf{v}_{\theta}(t,m,e,d)$
@@ C.1 Conditional Flow Matching
      for each minibatch of samples $(m_{0},m_{1})$ do
@@
            $m_{0}\sim\text{prior distribution}$
            $m_{t}\leftarrow t\cdot m_{1}+(1-t)\cdot m_{0}$
            Compute the target velocity: $u_{t}\leftarrow m_{1}-m_{0}$
            Predict the velocity: $v_{t}\leftarrow\mathbf{v}(t,m_{t},e,d)$
            Compute the loss: $\mathcal{L}(\theta)\leftarrow\mathbb{E}\left[\|v_{t}-u_{t}\|^{2}\right]$
@@
Algorithm 1 Conditional Flow Matching Training Algorithm
Algorithm 2 presents the inference procedure. After training, the model is used to define a deterministic flow by solving an ordinary differential equation (ODE) starting from a sample from the prior distribution. The terminal state of this ODE corresponds to a sample from the conditional distribution given the observations and experimental conditions.
Together, these two procedures enable the model to learn and sample from complex conditional distributions without relying on stochastic sampling or iterative optimization during inference.
Output: Generated parameters $m$
$x(t=0)\sim\text{prior distribution}$
$x(t=1)\leftarrow\text{Solution}\frac{dx}{dt}=\mathbf{v}_{\theta}(t,x_{t},e,d)$
return $x(t=1)$
Algorithm 2 Conditional Flow Matching Inference Algorithm